# 1 Admission by Purchase of Interest

The capital accounts of the Maniquiz and Monte Partnership on Sept 30, 2018 were:

Maniquiz Capital (75% Profit Percentage) P 140,000

Monte, Capital (25% profit percentage) 56,000

Total P 196,000

On Oct 1. Galang was admitted to a 35% interest in the partnership when he purchased 35% of each existing partner’s capital for P 100,000, paid directly to Maniquez and Monte.


Required:

Determine the capital balances of Maniquiz, Monte and Galang after Galang’s admission to the partnership.


1
Expert's answer
2021-10-11T16:09:16-0400

"140000*0.35=49000"

"56000*0.35=19600"

"140000-49000=91000" Maniquiz capital "\frac{91000}{196000}=0.46=46\%" profit

"56000-19600=36400" Monte capital 19% profit

68600 Galang capital 35% profit